Finally, we have curtains around the SPEII to seal the microscope space from the rest of the lab. This way we now can work normally in the wetlab again with the lights on.
After a long wait over the holiday period, the repair of the EL6000 fluorescence unit from the SPE2 LEICA confocal has been completed. It can be used again for finding fluorescence using the eyepieces before imaging on the CLSM.
Image distortions due to vibrations on the system were reported last week. We elevated the vibration insulator silicon supports and temporarily solved the problem. New insulators will be placed during the next Leica service. Please notify us when stripe patterns appear in your images.
We have an old building that is leaky, a serious water spill occurred when the last storm hit the east wing of the Androclus building. The equipment (the SPEII) was not damaged due to the swift action of Martijn Moolenaar (thanks again). Apparently, the rain leaked into the building at multiple departments and it is a major concern of the maintenance department. Although patch-type preventive measures are in progress, recurrence of the flooding cannot be ruled out. We urge users to notify CCI staf asap when it occurs again.
To minimize damage due to such a recurring accident we moved the SPEII to the east wall of the room and lifted all wiring from the floor. The machine is in full operation again.
Clearly, another save location for the SPEII in the tightly packed CCI is high on our agenda.
A StorSimple backup system has been implemented by ITS. We have tested it for a couple of months and it seems to perform well. The extra network drive is available on all imaging machines (except for the Zeiss NLO system, that will be connected to the net in the coming weeks) and can be found at O:\Research\CCI\#equipment\##user##. Writing on the drive can only be done on our machines, users will be able to read only from these drives so data will be available but secured. We are implementing a secure logon system for users to read data from this drive.
The new 40x objective (HC PL APO CS2 NA1.3, oil immersion) has been equipped with the correct condenser prisms some time ago, but it was not mentioned here. So the DIC and fluorescence is fully upgraded on the machine. Please realize the greater sensitivity of the new objective. Intensity measurements acquired before June with the 40x will have lower values than the current 40x.
A Leica service engineer has visited us and repaired the DIC slider bug and updated driver software. Keep us posted when bugs appear.
We experience a problem with the DIC slider in the SPE-II microscope. Leica has been contacted and will be repairing it in the coming days.
We have setup the SPE2 confocal to room O.101 as was announced before. The system is functional again and the PC software has been updated. We will evaluate whether it has solved the many problems we face lately with the performance of the system.
This week we will move the Leica SPE2 to labspace O.101 to prepare for the maintenance on the aircos next week. We will try to minimize inconvenience for users by clustering the restructuring of CCI spaces for the new scopes that will arrive. The cryostate has been moved to the lab O.102. Watch this space for progress of the change at the CCI.